solve the equation by the method of factoring: 4x^3+6x^2=70x
Answer by greenestamps(13198)



4x^3+6x^2 = 70x

Get all terms on one side, with "0" on the other side:

4x^3+6x^2-70x = 0

Divide by the constant factor 2:

2x^3+3x^2-35x = 0

Factor out the common power of the variable:

x(2x^2+3x-35) = 0

Factor the remaining quadratic factor:

x(2x-7)(x+5) = 0

The solutions are when any of the factors on the left are zero:

x = 0
2x-7 = 0 --> x = 7/2
x+5 = 0 --> x = -5

Answer: The roots are -5, 0, and 7/2.